OpenStudy (anonymous):

Hi everyone : hope ur well so I have this log plus absolute question so. For ln (x-2) = ln (2-x) I know there is no solution however for ln |x-2| = ln |2-x| There is , someone can kindly how is that possible and also help me solve this thank u

OpenStudy (perl):

I don't think there is a solution in either case.

OpenStudy (anonymous):

\[\ln|x-2|=\ln\left(|-1||2-x|\right)=\ln|2-x|\]

OpenStudy (anonymous):

You can think of the absolute value terms \(|2-x|\) and \(|x-2|\) in terms of distance. The distance between \(x\) and \(2\) is always the same. \(|a-b|\) can be read as "the distance from \(a\) to \(b\)", and \(|b-a|\) as "the distance from \(b\) to \(a\)". You can expect the distance to be the same regardless of the starting point.